Stephanie's age is three years more than twice Bill's age. Stephanie is nine years older than Bill. How old is Stephanie? How old is Bill?
Answer by Maths68(1474)

Stephanie's age is three years more than twice Bill's age. Stephanie is nine years older than Bill. How old is Stephanie? How old is Bill?
Let
Present age of Stephanie = s
Present age of Bill = b
Given
Stephanie's age is three years more than twice Bill's age
s=3+2b...........(1)
Stephanie is nine years older than Bill
s=9+b............(2)
Plug in the value of s in (2)
3+2b=9+b
2b-b=9-3
b=6
Put the value of b in (2)
s=9+b............(2)
s=9+6
s=15


Present age of Stephanie = s = 15 years old
Present age of Bill = b = 6 years old